OpenStudy (anonymous):

the displacement (in centimeters) of a particle moving back and forth along a straight line is given by the equation of motion s = 2 sin πt + 2 cos πt, where t is measured in seconds. (Round your answers to two decimal places.) Find the avg velocity on the intervals (a) Find the average velocity during each time period? i) 1,2 ii) 1,1.1

OpenStudy (dumbcow):

velocity function is derivative of motion equation: to find avg velocity for interval [a,b] \[\rightarrow \frac{1}{b-a}\int\limits_{a}^{b} s'(t) dt\] which equals \[\frac{s(b) - s(a)}{b-a}\]

OpenStudy (anonymous):

ok i got the derivative of 2pi(cospit-sinpit) I substituted the b(2) and (a). I got -103.17 which is still wrong.